# Principal component analysis balancing prediction and approximation accuracy for spatial data

### v2

@@ Abstract
indicates joint senior authors
@@ 2.1 Classical and Predictive PCA
for each $l=1,\ldots,r$. Here, $Z=[X\ B]$, $B\in\mathbb{R}^{n\times m}$ contains $m$ thin-plate regression spline basis functions capturing the spatial effects, and in contrast to (1), normalization is done on the term $Z{\alpha_{l}}$ instead of $v_{l}$. As before, $Y^{(l)}$ is the residual after approximation, i.e. $Y^{(l)}=Y^{(l-1)}-Y^{(l-1)}\frac{\tilde{v}_{l-1}\tilde{v}_{l-1}^{\top}}{\lVert\tilde{v}_{l-1}\rVert_{2}}$ for $2\leq l\leq r$ and $Y^{(1)}=Y$. Note that the quantity $Y^{(l-1)}\tilde{v}_{l-1}$ corresponds to $\tilde{u}_{l-1}$ in the expression of PCA. Compared with (1), this formulation restricts each PC score $u_{l}=\frac{Z\alpha_{l}}{\lVert Z\alpha_{l}\rVert_{2}}$ to fall within the space spanned by the columns of $Z$. This ensures the spatial smoothness and the inclusion of information from covariates $X$ in the PC score $u_{l}$, and thus enforces predictability.
@@ 2.2 Evaluation Metrics for Dimension Reduction
3 Representative and Predictive PCA (RapPCA)
@@
$\min_{u,v,\alpha,\beta}\ f_{\gamma,\lambda_{1},\lambda_{2}}(u,v,\alpha,\beta):=\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}+\gamma\left\lVert u-(K\alpha+B\beta)\right\rVert_{2}^{2}+\lambda_{1}\alpha^{\top}K\alpha+\lambda_{2}\beta^{\top}Q\beta$
@@
As in Section 2, $Y^{(l)}$ denotes the residual $Y^{(l-1)}-\tilde{u}_{l-1}{\tilde{v}}_{l-1}^{\top}$ for $2\leq l\leq r$, and is set to the original data $Y$ when $l=1$. Moreover, $K{\in\mathbb{R}^{n\times n}}$ is a kernel matrix such that $K_{ij}=k(X_{i\cdot},X_{j\cdot})$ for some kernel function $k(\cdot,\cdot)$, with $X_{i\cdot}$ being the $i$th observation (row) of $X$. The kernel function $k(\cdot,\cdot)$ maps the covariates $X$ into a higher-dimensional space, and the kernel matrix $K$ encodes the pairwise similarities between each entry $X_{i\cdot}$ in this transformed space. Linear operations in the high-dimensional space effectively captures non-linear relationships in the original input space of $X$. This enables more flexibility than most existing predictive variants of PCA which specify linear relationships (e.g., Jandarov et al., 2017; Bose et al., 2018; Vu et al., 2020; Shang and Zhou, 2022). Common choices of kernel functions include linear kernel, $k(x,x^{\prime})=x^{\top}x^{\prime}$, and Gaussian kernel, $k_{h}(x,x^{\prime})=\exp\left(-h\lVert x-x^{\prime}\rVert^{2}\right)$ (Hofmann et al., 2008; Murphy, 2012). The columns of $B{\in\mathbb{R}^{n\times m}}$ are evaluations of $m$ spline basis functions, e.g., thin-plate regression splines as in predictive PCA, at the spatial coordinates $\{(s_{i1},s_{i2})\}_{i=1}^{n}$, and $Q{\in\mathbb{R}^{m\times m}}$ is the penalty matrix induced by the choice of smoothing splines. For example, the penalty $Q$ corresponding to the thin-plate regression splines is the wiggliness penalty defined in Wood (2003); for P-splines, the penalty term depends on the differences between consecutive entries of $\beta$, i.e., $\sum_{k}(\beta_{k}-\beta_{k-1})^{2}$ (Wood, 2017).
@@
In (3), the first term $\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}$ measures the approximation gap of the score $u$ and loading $v$, similar to the objective of minimization for classical PCA. The second term reflects the distance between the score $u$ and the model space specified by $K$ and $B$, and encourages $u$ to be predictable based on the covariates and spatial coordinates. The constraint $u=Y^{(l)}v$ ensures a similar relationship between the PC scores and loadings as in classical PCA.
@@
The computation of the optimization problem in (3) could suffer when $K$ or $Q$ is near-singular. Therefore in practice, we replace $K$ and $Q$ with $\tilde{K}:=K+\delta I_{n}$ and $\tilde{Q}:=Q+\delta I_{m}$ for a small, data-independent constant $\delta$, and identity matrices $I_{n}$ and $I_{m}$ in the penalty terms to avoid near-singularity; see also Li et al. (2019) for similar handling. This leads to
[…]
@@
Here, $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ are tuning parameters, among which $\gamma$ is of most interest as it controls the trade-off between predictability and representability, with larger $\gamma$ imposing higher penalty on unpredictable scores $u:=Y^{(l)}v$. The third and fourth terms in (4) are $\ell_{2}$ penalties that enforce the identifiability of the model parameters and help to avoid overfitting the data.
@@ 21      Calculate the $l$th PC score: $\tilde{u}_{l}\leftarrow Y^{(l)}\tilde{v}_{l}$;
The computational complexity of the procedure described above is $O\left((n+m)^{3}+np^{2}\right)$, where we recall that $n$ is the sample size, $p$ is the dimensionality of $Y$, and $m$ is the number of spline basis functions included in $B$. This algorithm is computationally simple in that it obtains the optimal solution with an explicit expression, as opposed to iterative numerical optimization procedures. This algorithm depends on the specification of several quantities. For the hyper-parameters $\gamma^{(l)},\lambda_{1}^{(l)},\lambda_{2}^{(l)}$, one could adopt cross-validation and choose the combination that minimizes TMSE or other metrics (e.g., MSPE/MSRE) depending on the targeted use case, sequentially for each PC. Taking TMSE as an example, for the $l$th PC and for each combination of $\gamma^{(l)},\lambda_{1}^{(l)},\lambda_{2}^{(l)}$, we calculate $\tilde{u}_{l},\tilde{v}_{l}$ on the training set, predict $\hat{u}_{\text{tst}}$ on the validation set, and then examine $\text{TMSE}_{l}=n_{\text{tst}}^{-1}\left\lVert Y^{(l)}_{\text{tst}}-\hat{u}_{\text{tst}}\tilde{v}_{l}^{\top}\right\rVert_{2}^{2}$ to choose the optimal combination minimizing $\text{TMSE}_{l}$. Specification of the dimension $m$ of $B$ is not critical due to the inclusion of penalty on $\beta$ (Wood, 2017). One can start with a large $m$ (e.g., close to the sample size $n$) and the degree of freedom of $\beta$ will be controlled data-adaptively via $\lambda_{2}$.
@@ 4 Simulations
We examine metrics including TMSE, MSPE and MSRE-trn for 100 replicates of data, in terms of their overall magnitudes, as well as the breakdown by each PC and/or $\gamma$ (the tuning parameter). We also report the individual prediction MSEs for each PC, which is $\text{MSE}_{l}=n_{\text{tst}}^{-1}\lVert\hat{u}_{l}-u^{*}_{l}\rVert_{2}^{2}$, recalling from Section 2.2 that $u^{*}_{l}=Y_{\text{tst}}\tilde{v}_{l}$. We extract $r=3$ PCs sequentially with each of the three candidate algorithms, PCA, predictive PCA and RapPCA. The optimal combination of $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ is selected by minimizing TMSE via 10-fold cross-validation. For each PC score $\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3}$, test set predictions are obtained via a two-step procedure (see, e.g., Wai et al., 2020), where we first train a random forest model, and then conduct spatial smoothing on the residuals with thin-plate regression splines (TPRS) (Wood, 2003). We discuss the role played by the primary tuning parameter $\gamma$ in this section. Figure S3 for RapPCA in Appendix B.2 provides further details about the impact of $\lambda_{1},\lambda_{2}$ on the solution.
@@ 4.1 Scenario 1: PCs with Equal Contribution
Comparing the two existing methods, it is expected that in this scenario predictive PCA would outperform classical PCA in prediction without severely compromising approximation accuracy, since the predictable PCs explain a similar amount of variability compared to the unpredictable ones. Consequently, predictive PCA would achieve better overall performance as reflected by TMSE. RapPCA flexibly interpolates between classical and predictive PCA, and is expected to have comparable or better performance than predictive PCA.
Panel A in Figure 1 presents the breakdown of MSPE, MSRE-trn and TMSE for the first PC by $\gamma$, in comparison to PCA and predictive PCA for this scenario. We only present this breakdown for the first PC but not the subsequent ones because starting from the second PC, different PCA algorithms will have different residuals $Y^{(l)}$ (recall Equations 12 and 4), and the PC-specific breakdown of metrics is consequently no longer comparable. We observe that MSPE decreases and training MSRE increases as we increase $\gamma$ for RapPCA, since this imposes a greater penalty on prediction errors in optimization.
@@ 5.1 Analysis of Seattle Traffic-Related Air Pollution Data
We are interested in extrapolating the pollutant measurements to unobserved locations in the same study region. Because of the autocorrelation between these measurements, predicting each of them separately might not always lead to sensible results; instead, dimension reduction is needed before building spatial prediction models. Dimension reduction also allows us to understand the health effects of pollutant mixtures rather than separate pollutants alone. As in Section 4, we compare the performance of RapPCA with classical and predictive PCA, each extracting the top 3 PCs from the 23 measurements of air pollutants. Due to the high dimensionality of geographical covariates, we first ran PCA on these 189 covariates and used the top 15 PCs as predictors for predictive PCA. PCA and RapPCA were applied without this additional step. We examine the overall metrics as well as the individual prediction MSEs for each PC via 10-fold cross-validation following the same spatial prediction approach as Section 4, namely, a random forest model followed by spatial smoothing via TPRS on the residuals (Wai et al., 2020).
@@
Table 1: Comparison of overall metrics and individual prediction MSEs for each PC, assessed by 10-fold cross-validation on the Seattle traffic-related air pollution data
@@ 5.2 Analysis of HER2-Positive Breast Tumor Spatial Transcriptomics
The first application is more similar to the analysis in Section 5.1, where we are interested in predicting the gene expression PC scores at tissue locations with no measurements. This prediction problem is involved, for example, in the reconstruction of high-resolution spatial maps of gene expression (Gryglewski et al., 2018; Shang and Zhou, 2022). For each dimension reduction method, we extract the top 20 PCs from the expression of all genes, and conduct 10-fold cross-validation to assess TMSE, MSPE and MSRE-trn, where spatial prediction is done by random forest followed by TPRS smoothing as in previous sections. For spatial PCA, we also include results from its built-in spatial prediction function (using the SpatialPCA R package) for completeness. Table 2 summarizes the overall metrics along with individual prediction MSEs for the top 3 PCs. RapPCA demonstrates significant advantage in overall prediction accuracy (MSPE) compared to all other methods, followed by predictive PCA which also has the lowest per-PC MSEs for the top 3 PCs. Also, consistent with findings from previous sections, RapPCA achieves the optimal trade-off between prediction and approximation gaps, as reflected by TMSE.
@@
PCA 272.21 32.24 239.97 6.64 5.09 3.18
PredPCA 278.08 31.85 246.15 3.93 4.12 2.26
SpatialPCA 292.78 46.09 246.70 6.84 5.89 3.69
SpatialPCA: built-in 296.59 49.90 246.70 6.81 5.96 3.81
RapPCA 271.35 20.56 250.80 5.59 4.86 3.79
Table 2: Comparison of overall metrics and individual prediction MSEs for each of the top 3 PCs, assessed by 10-fold cross-validation on the breast tumor data

### v3

@@ Abstract
indicates joint senior authors[funding information]. All authors declare no conflicts of interest.
@@ 2.1 Classical and Predictive PCA
for each $l=1,\ldots,r$. Here, $Z=[X\ B]$, $B\in\mathbb{R}^{n\times m}$ contains $m$ thin-plate regression spline (TPRS; Wood, 2003) basis functions capturing the spatial effects, and in contrast to (1), normalization is done on the term $Z{\alpha_{l}}$ instead of $v_{l}$. As before, $Y^{(l)}$ is the residual after approximation, i.e. $Y^{(l)}=Y^{(l-1)}-Y^{(l-1)}\frac{\tilde{v}_{l-1}\tilde{v}_{l-1}^{\top}}{\lVert\tilde{v}_{l-1}\rVert_{2}}$ for $2\leq l\leq r$ and $Y^{(1)}=Y$. Note that the quantity $Y^{(l-1)}\tilde{v}_{l-1}$ corresponds to $\tilde{u}_{l-1}$ in the expression of PCA. Compared with (1), this formulation restricts each PC score $u_{l}=\frac{Z\alpha_{l}}{\lVert Z\alpha_{l}\rVert_{2}}$ to fall within the space spanned by the columns of $Z$. This ensures the spatial smoothness and the inclusion of information from covariates $X$ in the PC score $u_{l}$, and thus enforces predictability.
@@ 2.2 Evaluation Metrics for Dimension Reduction
3 The Proposed Method
3.1 Representative and Predictive PCA (RapPCA)
@@
$\min_{v,\alpha,\beta}\ f_{\gamma,\lambda_{1},\lambda_{2}}(v,\alpha,\beta):=\left\lVert Y^{(l)}-Y^{(l)}vv^{\top}\right\rVert_{F}^{2}+\gamma\left\lVert Y^{(l)}v-(K\alpha+B\beta)\right\rVert_{2}^{2}+\lambda_{1}\alpha^{\top}K\alpha+\lambda_{2}\beta^{\top}Q\beta$
@@
As in Section 2, $Y^{(l)}$ denotes the residual $Y^{(l-1)}-\tilde{u}_{l-1}{\tilde{v}}_{l-1}^{\top}$ for $2\leq l\leq r$, where $\tilde{u}_{l-1}=Y^{(l-1)}\tilde{v}_{(l-1)}$, and $Y^{(l)}$ is set to the original data $Y$ when $l=1$. Moreover, $K{\in\mathbb{R}^{n\times n}}$ is a kernel matrix such that $K_{ij}=k(X_{i\cdot},X_{j\cdot})$ for some kernel function $k(\cdot,\cdot)$, with $X_{i\cdot}$ being the $i$th observation (row) of $X$. The kernel function $k(\cdot,\cdot)$ maps the covariates $X$ into a higher-dimensional space, and the kernel matrix $K$ encodes the pairwise similarities between each entry $X_{i\cdot}$ in this transformed space. Linear operations in the high-dimensional space effectively captures non-linear relationships in the original input space of $X$. This enables more flexibility than most existing predictive variants of PCA which specify linear relationships (e.g., Jandarov et al., 2017; Bose et al., 2018; Vu et al., 2020; Shang and Zhou, 2022). Common choices of kernel functions include linear kernel, $k(x,x^{\prime})=x^{\top}x^{\prime}$ and quadratic kernel $k(x,x^{\prime})=(1+x^{\top}x^{\prime})^{2}$. Other choices, such as Gaussian kernel, $k_{h}(x,x^{\prime})=\exp\left(-h\lVert x-x^{\prime}\rVert^{2}\right)$, or higher-order polynomial kernel $k_{h}(x,x^{\prime})=(1+x^{\top}x^{\prime})^{h}$ may include an additional tuning parameter, $h$ (Hofmann et al., 2008; Murphy, 2012). The columns of $B{\in\mathbb{R}^{n\times m}}$ are evaluations of $m$ spline basis functions, e.g., thin-plate regression splines as in predictive PCA, at the spatial coordinates $\{(s_{i1},s_{i2})\}_{i=1}^{n}$, and $Q{\in\mathbb{R}^{m\times m}}$ is the penalty matrix induced by the choice of smoothing splines. For example, the penalty $Q$ corresponding to thin-plate regression splines is the wiggliness penalty defined in Wood (2003); for P-splines, the penalty term depends on the differences between consecutive entries of $\beta$, i.e., $\sum_{k}(\beta_{k}-\beta_{k-1})^{2}$ (Wood, 2017).
Problem (3) is equivalent to optimizing
$\tilde{f}_{\gamma,\lambda_{1},\lambda_{2}}(u,v,\alpha,\beta):=\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}+\gamma\left\lVert u-(K\alpha+B\beta)\right\rVert_{2}^{2}+\lambda_{1}\alpha^{\top}K\alpha+\lambda_{2}\beta^{\top}Q\beta,$
@@
subject to an additional constraint $u=Y^{(l)}v$. Its first term $\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}$ measures the approximation gap of the score $u$ and loading $v$, similar to the objective of minimization for classical PCA. The second term reflects the distance between the score $u$ and the model space specified by $K$ and $B$, and encourages $u$ to be predictable based on the covariates and spatial coordinates. The implicit constraint $u=Y^{(l)}v$ ensures a similar relationship between the PC scores and loadings as in classical PCA.
@@
The computation of the optimization problem in (3) could suffer when $K$ or $Q$ is near-singular. Therefore, in practice, we replace $K$ and $Q$ in the penalty term with $\tilde{K}:=K+\delta I_{n}$ and $\tilde{Q}:=Q+\delta I_{m}$ for a small, data-independent constant $\delta$, and identity matrices $I_{n}$ and $I_{m}$, to avoid near-singularity; see also Li et al. (2019) for similar handling.
@@
Among the tuning parameters $\gamma$, $\lambda_{1}$ and $\lambda_{2}$, $\gamma$ is of most interest as it controls the trade-off between predictability and representability, with larger $\gamma$ imposing higher penalty on unpredictable scores $u=Y^{(l)}v$. The third and fourth terms in (3) are $\ell_{2}$ penalties that enforce the identifiability of the model parameters and help to avoid overfitting the data.
@@ 21      Calculate the $l$th PC score: $\tilde{u}_{l}\leftarrow Y^{(l)}\tilde{v}_{l}$;
3.2 Computational Considerations
When $m$ spline basis functions are included in $B$ to represent a $p$-dimensional $Y$ observed over $n$ samples, the computational complexity of the procedure described in Algorithm 1 is $O\left((n+m)^{3}+np^{2}\right)$. This algorithm is computationally simple in that it obtains the optimal solution with an explicit expression, as opposed to iterative numerical optimization procedures.
The proposed algorithm depends on the specification of several quantities. For the hyper-parameters $\gamma^{(l)},\lambda_{1}^{(l)},\lambda_{2}^{(l)}$, one could adopt cross-validation and choose the combination that minimizes TMSE or other metrics (e.g., MSPE/MSRE) depending on the targeted use case, sequentially for each PC. Taking TMSE as an example, for the $l$th PC and for each combination of $\gamma^{(l)},\lambda_{1}^{(l)},\lambda_{2}^{(l)}$, we calculate $\tilde{u}_{l},\tilde{v}_{l}$ on the training set, predict $\hat{u}_{\text{tst}}$ on the validation set, and then examine $\text{TMSE}_{l}=n_{\text{tst}}^{-1}\left\lVert Y^{(l)}_{\text{tst}}-\hat{u}_{\text{tst}}\tilde{v}_{l}^{\top}\right\rVert_{2}^{2}$ to choose the optimal combination minimizing $\text{TMSE}_{l}$. When necessary, the additional tuning parameter for kernel function, such as $h$ for the Gaussian kernel, can be chosen together with $\gamma,\lambda_{1},\lambda_{2}$ using cross-validation. Specification of the dimension $m$ of $B$ is not critical due to the penalty on $\beta$ (Wood, 2017). One can start with a large $m$ (e.g., close to the sample size $n$) and the degree of freedom of $\beta$ will be controlled data-adaptively via $\lambda_{2}$.
The implementation of Algorithm 1 also depends on the number $r$ of PCs. Since increasing $r$ will reduce the total representation error and increase the total prediction error, the selection can be done by examining the trajectories of MSPE, or cumulative sum of PC-wise prediction errors, along with MSRE with respect to $r$, and identifying the ‘elbow’ of these curves. This reflects the threshold beyond which adding more PCs would not alter either the prediction or representation meaningfully. We will present an illustrative example (Figure 5) for this procedure in Section 5.2, which involves a high-dimensional breast tumor dataset.
@@ 4 Simulations
We examine metrics including TMSE, MSPE and MSRE-trn for 100 replicates of data, in terms of their overall magnitudes, as well as the breakdown by each PC and/or $\gamma$ (the tuning parameter). We also report the individual prediction MSEs for each PC, which is $\text{MSE}_{l}=n_{\text{tst}}^{-1}\lVert\hat{u}_{l}-u^{*}_{l}\rVert_{2}^{2}$, recalling from Section 2.2 that $u^{*}_{l}=Y_{\text{tst}}\tilde{v}_{l}$. We extract $r=3$ PCs sequentially with each of the three candidate algorithms, PCA, predictive PCA and RapPCA. The optimal combination of $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ is selected by minimizing TMSE via 10-fold cross-validation on a grid where $\gamma,\lambda_{1}$ and the ratio $\lambda_{2}/\lambda_{1}$ each independently varies from 0.05 to 5 with varying spacings [^1] We use 0.05 for hyperparameter values up to 0.1; 0.1 for hyperparameter values between 0.1 and 1; 1 for hyperparameter values over 1., plus the combination $\gamma=\lambda_{1}=\lambda_{2}=0$ representing classical PCA as a special case. We illustrate the performance of RapPCA with polynomial kernel $k_{h}(x,x^{\prime})=(1+x^{\top}x^{\prime})^{h}$, with $h=1$ (linear) for scenarios 1 and 2, and $h=2$ for scenario 3. We adopt thin-plate regression splines implemented by the mgcv R package to capture the spatial pattern in each PC, with dimension $m$ equal to the size of training data for RapPCA, and the penalty term is chosen as cubic spline penalty with second order derivatives. Since predictive PCA does not directly handle high-dimensionality, we truncate the spline basis terms $B$ at $m=10$ for predictive PCA. The value of $\delta$ is fixed at 0.05, which is small relative to $\gamma,\lambda_{1},\lambda_{2}$ but helps avoid near-singular penalties. For each PC score $\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3}$, test set predictions are obtained via a two-step procedure (see, e.g., Wai et al., 2020), where we first train a random forest model, and then conduct spatial smoothing on the residuals with thin-plate regression splines. We discuss the role played by the primary tuning parameter $\gamma$ in this section. Figure S3 in Appendix B.2 provides further details about the impact of $\lambda_{1},\lambda_{2}$ on the solution of RapPCA.
@@ 4.1 Scenario 1: PCs with Equal Contribution
Comparing the two existing methods, it is expected that in this scenario predictive PCA would outperform classical PCA in prediction without severely compromising approximation accuracy, as the predictable PCs explain a similar amount of variability compared to the unpredictable ones. Consequently, predictive PCA would achieve better overall performance as reflected by TMSE. RapPCA flexibly interpolates between classical and predictive PCA, and is expected to have comparable or better performance than predictive PCA.
Panel A in Figure 1 presents the breakdown of MSPE, MSRE-trn and TMSE for the first PC by $\gamma$, in comparison to PCA and predictive PCA for this scenario. We only present this breakdown for the first PC but not the subsequent ones because starting from the second PC, different PCA algorithms will have different residuals $Y^{(l)}$ (recall Equations 12 and 3), and the PC-specific breakdown of metrics is consequently no longer comparable. We observe that MSPE decreases and training MSRE increases as we increase $\gamma$ for RapPCA, since this imposes a greater penalty on prediction errors in optimization.
@@ 5.1 Analysis of Seattle Traffic-Related Air Pollution Data
We are interested in extrapolating the pollutant measurements to unobserved locations in the same study region. Because of the autocorrelation between these measurements, predicting each of them separately might not always lead to sensible results; instead, dimension reduction is needed before building spatial prediction models. Dimension reduction also allows us to understand the health effects of pollutant mixtures rather than separate pollutants alone. As in Section 4, we compare the performance of RapPCA with classical and predictive PCA, each extracting the top 3 PCs from the 23 measurements of air pollutants. Due to the high dimensionality of geographical covariates, we first ran PCA on these 189 covariates and used the top 15 PCs as predictors for predictive PCA. PCA and RapPCA were applied without this additional step. We chose a linear kernel and the same configurations for thin-plate regression splines as in Section 4. We adopted the same grid for the tuning of hyperparameters ($\lambda_{1},\lambda_{2}/\lambda_{1}$) from 0.05 to 5 and the same value $\delta=0.05$, but expanded the candidate range of $\gamma$ to $[0.05,50]$ to accommodate wider ranges of prediction and representation errors. We examine the overall metrics as well as the individual prediction MSEs for each PC via 10-fold cross-validation following the same spatial prediction approach as Section 4, namely, a random forest model followed by spatial smoothing via TPRS on the residuals (Wai et al., 2020).
@@
Table 1: Comparison of overall metrics and individual prediction MSEs for each PC, assessed by 10-fold cross-validation on the Seattle traffic-related air pollution data. Standard deviations of these metrics evaluated across 10 cross-validation folds are reported in Table S1 in Appendix B.3.
@@ 5.2 Analysis of HER2-Positive Breast Tumor Spatial Transcriptomics
The first application is more similar to the analysis in Section 5.1, where we are interested in predicting the gene expression PC scores at tissue locations with no measurements. This prediction problem is involved, for example, in the reconstruction of high-resolution spatial maps of gene expression (Gryglewski et al., 2018; Shang and Zhou, 2022). For each dimension reduction method, we extract the top 20 PCs from the expression of all genes, and conduct 10-fold cross-validation to assess TMSE, MSPE and MSRE-trn, where spatial prediction is done by random forest followed by TPRS smoothing as in previous sections. For spatial PCA, we also include results from its built-in spatial prediction function (using the SpatialPCA R package) for completeness. There are no covariates involved and we only use spatial information for both the spatial extrapolation and domain detection tasks. We adopt the same spatial splines and hyperparameter ranges as in the air pollution study in Section 5.1, and recall that the dimension $m$ of spatial basis is truncated to 10 for predictive PCA whereas RapPCA sets this dimension to the size of the training data. Table 2 summarizes the overall metrics along with individual prediction MSEs for the top 3 PCs. RapPCA demonstrates significant advantage in overall prediction accuracy (MSPE) compared to all other methods, followed by predictive PCA which also has the lowest per-PC MSEs for the top 3 PCs. Consistent with findings from previous sections, RapPCA achieves the optimal trade-off between prediction and approximation gaps, as reflected by TMSE.
@@
PCA 272.21 32.24 219.97 6.64 5.09 3.18
PredPCA 278.08 31.85 241.91 3.93 4.12 2.26
SpatialPCA 292.78 46.09 236.88 6.84 5.89 3.69
SpatialPCA: built-in 296.59 49.90 236.88 6.81 5.96 3.81
RapPCA 271.35 20.56 243.27 5.59 4.86 3.79
Table 2: Comparison of overall metrics and individual prediction MSEs for each of the top 3 PCs, assessed by 10-fold cross-validation on the breast tumor data. Standard deviations of these metrics evaluated across 10 cross-validation folds are reported in Table S2 in Appendix B.3.
@@
To provide some heuristics on our choice of $r$ as described in Section 3.2, Figure 5 visualizes the trends for the cumulative sum of PC-wise prediction MSEs, along with MSRE-trn of RapPCA. We observe that the performance of RapPCA starts to stabilize for $r$ between 10 and 15, justifying $r=20$ to be sufficient.
Figure 5: Changes in cumulative prediction MSE and MSRE-trn with the number of PCs for RapPCA on the breast tumor dataset.